OpenStudy (anonymous):

Use the Pythagorean Theorem to find the length of the missing side of a right triangle whose legs equal √2 and 2√15.

OpenStudy (adamk):

a^2 + b^2 = c^2 4 + 30 = c^2

OpenStudy (kc_kennylau):

@AdamK not 30 :)

OpenStudy (adamk):

Sorry the second term, b, should be 60

OpenStudy (adamk):

Oh, I got it :)

OpenStudy (adamk):

Take the square root of 64 and you have c

OpenStudy (anonymous):

il pipe you

OpenStudy (anonymous):

We'll the square root of 2 I rounded to 1 cuz it couldn't round any higher and the square root of 15 multiplied by 2 would be approximately 8 so 1^2+8^2= 65 or approximately 8-8.5^2 I think